# HPS: Hard Preference Sampling for Human Preference Alignment

### v1

@@ Abstract
Aligning Large Language Model (LLM) responses with human preferences is vital for building safe and controllable AI systems. While preference optimization methods based on Plackett-Luce (PL) and Bradley-Terry (BT) models have shown promise, they face challenges such as poor handling of harmful content, inefficient use of dispreferred responses, and, specifically for PL, high computational costs. To address these issues, we propose Hard Preference Sampling (HPS), a novel framework for robust and efficient human preference alignment. HPS introduces a training loss that prioritizes the most preferred response while rejecting all dispreferred and harmful ones. It emphasizes “hard” dispreferred responses—those closely resembling preferred ones—to enhance the model’s rejection capabilities. By leveraging a single-sample Monte Carlo sampling strategy, HPS reduces computational overhead while maintaining alignment quality. Theoretically, HPS improves sample efficiency over existing PL methods and maximizes the reward margin between preferred and dispreferred responses, ensuring clearer distinctions. Experiments on HH-RLHF and PKU-Safety datasets validate HPS’s effectiveness, achieving comparable BLEU and reward scores while greatly improving reward margins and thus reducing harmful content generation.
@@ 1 Introduction
Firstly, we introduce the HPS framework to enhance human preference alignment. Specifically, we first propose a training loss that fine-tunes LLMs to robustly prefer the most desired response while rejecting all dispreferred and potentially harmful ones. Moreover, HPS leverages insights from supervised, metric, and contrastive learning (Schroff et al., 2015; Oh Song et al., 2016; Robinson et al., 2020), emphasizing the importance of “hard” examples—dispreferred responses closely resembling the preferred ones. Accordingly, HPS develops a hard preference sampling strategy to prioritize such hard examples, enabling the model to distinguish between preferred and highly similar dispreferred responses more effectively. To ensure efficiency, HPS is then reformulated into a sampling approach, using a single Monte Carlo sampling to select a single dispreferred response per training iteration. This innovation significantly reduces computational overhead compared to PL which requires all dispreferred responses for each prompt.
@@ 2 Related Work
However, RLHF’s on-policy nature introduces challenges. It requires learning a reward model from data as a preliminary step, leading to a complex two-stage optimization process. Recent advancements in preference alignment techniques have sought to simplify this process by enabling direct alignment through a single loss function (Rafailov et al., 2024; Park et al., 2024; Azar et al., 2024; Wu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Ji et al., 2024a; Chen et al., 2024; Song et al., 2024). While these techniques streamline optimization, they face limitations such as poor handling of harmful content, inefficient utilization of dispreferred responses, and high computational costs.
To address these limitations, we propose HPS, a novel framework for robust and efficient human preference alignment. HPS prioritizes the most preferred response while explicitly rejecting dispreferred and harmful ones. By emphasizing “hard” dispreferred responses—those closely resembling preferred ones—it improves rejection capabilities. Additionally, a single-sample Monte Carlo strategy reduces computational overhead while maintaining strong alignment quality.
@@ 3 Preliminaries
For explicit methods, Reinforcement Learning from Human Feedback (RLHF) is widely used. RLHF trains a reward model $r_{\mathbf{{\boldsymbol{\theta}}}}$ to learn response rankings in $\mathcal{D}$, then fine-tunes LLM $\pi_{\text{SFT}}$ using policy-gradient algorithms like PPO (Schulman et al., 2017) to generate higher-preference responses. Refer to previous works (Ziegler et al., 2019; Ouyang et al., 2022) for further details.
@@
$\mathcal{L}_{j}(d)\!=\!-\log\!\Big{(}{e^{r_{{\boldsymbol{\theta}}}(x,y_{\tau(j)}}/\sum\nolimits_{k=j}^{n}e^{r_{{\boldsymbol{\theta}}}(x,y_{\tau(k)}})}\Big{)}.$ (5)
@@ 4.2 Hard Preference Sampling for Alignment
where the model is encouraged to rank $y_{\tau(1)}$ above all dispreferred and potentially harmful responses $\{y_{\tau(i)}\}_{i=2}^{n}$. We use the DPO implicit reward as mentioned in Eqn. 3 here.
However, this loss treats all dispreferred responses $\{y_{\tau(i)}\}_{i=2}^{n}$ equally, ignoring their varying levels of informativeness. Previous works in supervised, metric, and contrastive learning (Schroff et al., 2015; Oh Song et al., 2016; Robinson et al., 2020) demonstrate that “hard” examples—those closely resembling the correct output but still incorrect—are particularly useful for learning. In our context, hard dispreferred responses are those that are highly similar to $y_{\tau(1)}$ yet dispreferred or harmful. Training the model to distinguish $y_{\tau(1)}$ from the hardest dispreferred response $y_{\tau(2)}$ enables it to reject less preferred responses $\{y_{\tau(i)}\}_{i=3}^{n}$ more effectively. Thus, harder dispreferred responses should be penalized more heavily during training.
@@
where $r^{*}(x,y)$ is the inaccessible optimal reward model defined in Sec. 3 and can provide the ground-truth rewards, $p(y)$ is the probability density of dispreferred response $y$, and $Z$ is the partition function for normalization. For each ranked response $y_{\tau(i)}$, we can either directly access its reward $r_{\text{est}}$ if available in the dataset $\mathcal{D}$ or estimate it using a pretrained human preference-aligned reward model, $r_{\text{est}}(x,y_{\tau(i)})\approx r^{*}(x,y_{\tau(i)})$. Without loss of generality, we first formulate the Eqn. (6) in the expectation form:
@@
$q(x,y)={e^{\gamma\cdot r_{\text{est}}(x,y)}}/{\sum\nolimits_{i=2}^{n}e^{\gamma\cdot r_{\text{est}}(x,y)}}.$
@@
Reducing Training Costs with Flexible Sampling. Although this approach improves alignment, computing rewards for all $n$ responses and backpropagating through them can be computationally expensive. To address this, we propose to sample only a single importance-weighted dispreferred response $y\sim q(y)$ and incorporate it into the loss function Eqn. (9) for each prompt $x$ in practice, which works very well as shown in Sec. 6 and also significantly reduces computational and memory overhead. By focusing more on the hard dispreferred responses, our method retains robust alignment while greatly improving training efficiency.
@@ 5.1 Sample Complexity Analysis
where $\mathcal{L}_{{\boldsymbol{\theta}}}$ and $\mathcal{L}_{\text{PL}}$ respectively denote our HPS loss in Eqn. 9, and the PL loss in Eqn. 5. Then we pose necessary assumptions widely used in network and RLHF analysis (Zhu et al., 2023; Li et al., 2024; Ozay, 2019).
@@
Theorem 1 shows the bounded distance between the optimum solution $\boldsymbol{\theta}_{\text{HPS}}$ of our HPS loss and the ground-truth optimum $\boldsymbol{\theta}^{*}$ and indicates its good approximation.
Similarly, Theorem 1 also indicates the distance between the optimum solution $\boldsymbol{\theta}_{\text{PL}}$ of PL loss and the ground-truth $\boldsymbol{\theta}^{*}$ is bounded by $\Psi_{2}$. Now we compare the optimums of our HPS and PL by comparing the bounded distances $\Psi_{1}$ and $\Psi_{2}$ to the ground-truth $\boldsymbol{\theta}^{*}$. $\Psi_{1}$ represents an asymptotic error bound of $\mathcal{O}\big{(}\frac{n}{\sqrt{m}}\big{)}$, while $\Psi_{2}$ represents an asymptotic error bound of $\mathcal{O}\big{(}\frac{n^{2}}{\sqrt{m}}\big{)}$. This indicates that, given the same amount of training data, our HPS achieves better preference alignment performance compared to PL. Specifically, the optimum solution $\boldsymbol{\theta}_{\text{HPS}}$ derived from HPS loss is closer to the desired ground-truth $\boldsymbol{\theta}^{*}$ than the solution obtained from PL loss. This suggests that HPS improves sample efficiency, making it advantageous in scenarios with limited data or when faster convergence to the true parameter is desired.
@@ 5.2 Reward Margin Analysis
Furthermore, to examine the global minimizer of our HPS training loss $\mathcal{L}_{{\boldsymbol{\theta}}}$, we analyze the optima of the training loss $\mathcal{L}_{\Pi}^{*}$ in Eqn. 11, since we have proved their good approximation in Theorem 2. Without loss of generality, when the number of dispreferred response samples $N=n-1\rightarrow\infty$, we can remove the $\log{N}$ from the HPS training loss $\mathcal{L}_{{\boldsymbol{\theta}}}$ as it does not change the minimizers and the geometry of the loss surface, and obtain a limiting objective:
@@ 6 Experiments
Baselines. In experiments, we use a supervised fine-tuned LLaMA3-8B checkpoint (Dong et al., 2024; Dubey et al., 2024) as both baseline SFT and reference model. We test the SFT model and integrate our HPS into several preference optimization methods, including DPO (Rafailov et al., 2024) (BT&PL), EXO (Ji et al., 2024a) (BT&PL), IPO (Azar et al., 2024), SPPO (Wu et al., 2024), and NCA (Chen et al., 2024).
@@
Datasets. We use two popular datasets, HH-RLHF (Bai et al., 2022a) and PKU-SafeRLHF (Ji et al., 2024b), focusing on helpfulness and safety (Lambert et al., 2024; Fourrier et al., 2024). HH-RLHF is multi-turn, while PKU-SafeRLHF contains single question-answer pairs. Each prompt in datasets includes two responses with human-rated preferences. To test different methods, for each prompt, we adopt LLaMA3 (Dong et al., 2024) to generate 100 responses, and generate their rankings and rewards using a top-10 safety-ranked reward model (Liu et al., 2024) from the RewardBench benchmark (Lambert et al., 2024).
@@
Evaluation Metrics. We evaluate response quality and harmful content rejection. We use BLEU (Papineni et al., 2002) to assess the text quality by comparing responses to ground-truth preferred answers and adopt a strong reward model (Wang et al., 2024) to measure the level of human preference gained. We also compute reward margins (RMs) from various implicit reward models (Table 1) to quantify the gap between preferred and dispreferred responses, where higher RM scores indicate better preference alignment without harmful or biased outputs.
@@
Human evaluation remains the gold standard for assessing response quality, where annotators compare two responses per question to select the better one or declare a tie. Recent LLMs like Qwen $2.5$ (Yang et al., 2024) closely align with human preferences, validated by benchmarks like the Open LLM Leaderboard (Fourrier et al., 2024), chatbot-arena-leaderboard (Chiang et al., 2024), and RewardBench (Lambert et al., 2024). We use Qwen-2.5-Instruct to assess response quality and win rates. For evaluation robustness, we sample $N=5$ responses per method and report the highest-scoring one for each metric.
@@
Implementation. Due to computational constraints, we apply LoRA (Hu et al., 2021) for efficient fine-tuning with a rank of $8$ and scaling factor $\alpha=16$. The KL penalty strength $\beta$ is set to $0.1$, following DPO. We use a sequence length of $2,048$ tokens for both training and inference, with a sampling temperature of $0.9$ during inference. We set the number of responses $n\!=\!5$ for all PL-based methods. The GPU fine-tuning time for the PL-based, BT-based, and HPS-based methods is $168.4\!\pm\!1.9$ hours, $62.8\!\pm\!1.1$ hours, and $64.4\!\pm\!0.8$ hours, respectively. This demonstrates that our proposed HPS method can significantly improve efficiency compared to the PL-based method, achieving a $61.76\%$ reduction in fine-tuning time, and validates Theorem 1.

### v2

@@ Abstract
Aligning Large Language Model (LLM) responses with human preferences is vital for building safe and controllable AI systems. While preference optimization methods based on Plackett-Luce (PL) and Bradley-Terry (BT) models have shown promise, they face challenges such as poor handling of harmful content, inefficient use of dispreferred responses, and, specifically for PL, high computational costs. To address these issues, we propose Hard Preference Sampling (HPS), a novel framework for robust and efficient human preference alignment. HPS introduces a training loss that prioritizes the most preferred response while rejecting all dispreferred and harmful ones. It emphasizes “hard” dispreferred responses — those closely resembling preferred ones — to enhance the model’s rejection capabilities. By leveraging a single-sample Monte Carlo sampling strategy, HPS reduces computational overhead while maintaining alignment quality. Theoretically, HPS improves sample efficiency over existing PL methods and maximizes the reward margin between preferred and dispreferred responses, ensuring clearer distinctions. Experiments on HH-RLHF and PKU-Safety datasets validate HPS’s effectiveness, achieving comparable BLEU and reward scores while greatly improving reward margins and thus reducing harmful content generation. The source code is available at https://github.com/Yqcca/HPS.
@@ 1 Introduction
Firstly, we introduce the HPS framework to enhance human preference alignment. Specifically, we first propose a training loss that fine-tunes LLMs to robustly prefer the most desired response while rejecting all dispreferred and potentially harmful ones. Moreover, HPS leverages insights from supervised, metric, and contrastive learning (Schroff et al., 2015; Oh Song et al., 2016; Robinson et al., 2020), emphasizing the importance of “hard” examples—dispreferred responses closely resembling the preferred ones in the reward space (i.e. with close reward scores). Accordingly, HPS develops a hard preference sampling strategy to prioritize such hard examples, enabling the model to distinguish between preferred and highly similar dispreferred responses more effectively. To ensure efficiency, HPS is then reformulated into a sampling approach, using a single Monte Carlo sampling to select a single dispreferred response per training iteration. This innovation significantly reduces computational overhead compared to PL which requires all dispreferred responses for each prompt.
@@ 2 Related Work
However, RLHF’s on-policy nature introduces challenges. It requires learning a reward model from data as a preliminary step, leading to a complex two-stage optimization process. Recent advancements in preference alignment techniques have sought to simplify this process by enabling direct alignment through a single loss function (Rafailov et al., 2024; Park et al., 2024; Azar et al., 2024; Wu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Ji et al., 2024a; Chen et al., 2024). While these techniques streamline optimization, they face limitations such as poor handling of harmful content, inefficient utilization of dispreferred responses, and high computational costs.
In parallel, listwise preference learning methods (Song et al., 2024; Zhao et al., 2023; Liu et al., 2024b) offer a promising alternative. SLiC-HF (Zhao et al., 2023) is an alternative to RLHF-PPO (Schulman et al., 2017) by integrating the sequence-level contrastive method SLiC (Zhao et al., 2022) with human preference rankings. In LiPO-$\lambda$ (Liu et al., 2024b), it employs a listwise ranking objective with a Lambda weight, which assigns greater importance to response pairs with larger preference gaps. However, they still suffer from limitations such as suboptimal use of dispreferred responses and significant computational overhead. See Appendix A for details.
To address these limitations, we propose HPS, a novel framework for robust and efficient human preference alignment. HPS prioritizes the most preferred response while explicitly rejecting dispreferred and harmful ones. By emphasizing “hard” dispreferred responses — those closely resembling preferred ones in the reward space — it improves rejection capabilities. Additionally, a single-sample Monte Carlo strategy reduces computational overhead while maintaining strong alignment quality.
@@ 3 Preliminaries
For explicit methods, Reinforcement Learning from Human Feedback (RLHF) is widely used. RLHF trains a reward model $r_{\mathbf{{\boldsymbol{\theta}}}}$ to learn response rankings in $\mathcal{D}$, then fine-tunes LLM $\pi_{\text{SFT}}$ using policy-gradient algorithms like PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) to generate higher-preference responses. Refer to previous works (Ziegler et al., 2019; Ouyang et al., 2022) for further details.
@@
$\mathcal{L}_{j}(d)\!=\!-\log\!\Big{(}{e^{r_{{\boldsymbol{\theta}}}(x,y_{\tau(j)})}/\sum\nolimits_{k=j}^{n}e^{r_{{\boldsymbol{\theta}}}(x,y_{\tau(k)})}}\Big{)}.$ (5)
@@ 4.2 Hard Preference Sampling for Alignment
where the model is encouraged to rank $y_{\tau(1)}$ above all dispreferred and potentially harmful responses $\{y_{\tau(i)}\}_{i=2}^{n}$. We use the DPO implicit reward parameterization as mentioned in Eqn. 3 here. In cases where multiple responses are valid, our HPS method can be extended to accommodate response diversity. The details of this extension are provided in Appendix E.
However, this loss treats all dispreferred responses $\{y_{\tau(i)}\}_{i=2}^{n}$ equally, ignoring their varying levels of informativeness. Previous works in supervised, metric, and contrastive learning (Schroff et al., 2015; Oh Song et al., 2016; Robinson et al., 2020) demonstrate that “hard” examples — those closely resembling the correct output but still incorrect — are particularly useful for learning. In such settings, hard negatives are typically selected based on representation similarity to a positive anchor; however, in RLHF, where responses are generated autoregressively, obtaining effective sentence embeddings is impractical. Instead, in our context, hard dispreferred responses are those that are highly similar to $y_{\tau(1)}$ yet dispreferred or harmful in the reward space. Training the model to distinguish $y_{\tau(1)}$ from the hardest dispreferred response $y_{\tau(2)}$ enables it to reject less preferred responses $\{y_{\tau(i)}\}_{i=3}^{n}$ more effectively. Thus, harder dispreferred responses should be penalized more heavily during training.
@@
where $r^{*}(x,y)$ is the inaccessible optimal reward model defined in Sec. 3 and can provide the ground-truth rewards, $p(y)$ is the probability distribution of the dispreferred response $y$, and $Z$ is the partition function for normalization. For each ranked response $y_{\tau(i)}$, we can either directly access its reward $r_{\text{est}}$ if available in the dataset $\mathcal{D}$ or estimate it using a pretrained human preference-aligned reward model, $r_{\text{est}}(x,y_{\tau(i)})\approx r^{*}(x,y_{\tau(i)})$. Without loss of generality, we first formulate the Eqn. (6) in the expectation form:
@@
$q(x,y)={e^{\gamma\cdot r_{\text{est}}\left(x,y\right)}}/{\sum\nolimits_{i=2}^{n}e^{\gamma\cdot r_{\text{est}}\left(x,y_{\tau\left(i\right)}\right)}}.$ (9)
@@
Reducing Training Costs with Flexible Sampling. Although this approach improves alignment, computing rewards for all $n$ responses and backpropagating through them can be computationally expensive. To address this, we propose to sample only one dispreferred response $y\sim q(y)$ according to the importance-weighted distribution in Eqn. (9) given a prompt $x$. Thus, harder dispreferred responses will be sampled with higher probability and contribute more to the loss due to their higher $q(x,y)$. Then, we can incorporate the sampled dispreferred response $y$ into the loss function Eqn. (10) for each prompt $x$ in practice. This sampling technique works well as shown in Sec. 6 and also significantly reduces computational and memory overhead. By focusing more on the hard dispreferred responses, our method retains robust alignment while greatly improving training efficiency.
@@ 5.1 Sample Complexity Analysis
where $\mathcal{L}_{{\boldsymbol{\theta}}}$ and $\mathcal{L}_{\text{PL}}$ respectively denote our HPS loss in Eqn. 10, and the PL loss in Eqn. 5. Then we pose necessary assumptions widely used in network and RLHF analysis (Zhu et al., 2023; Li et al., 2024; Ozay, 2019).
@@
Theorem 1 shows the bounded distance $\Psi_{1}$ between the optimum solution $\boldsymbol{\theta}_{\text{HPS}}$ of our HPS loss and the ground-truth optimum $\boldsymbol{\theta}^{*}$ and indicates its good approximation. Theorem 1 also demonstrates the distance between the optimum solution $\boldsymbol{\theta}_{\text{PL}}$ of PL loss and the ground-truth $\boldsymbol{\theta}^{*}$ is bounded by $\Psi_{2}$. The difference in sample complexity bounds arises from the method used to sample dispreferred responses and the aggregation process for their associated scores given a prompt. Now we compare the optimums of our HPS and PL by comparing the bounded distances $\Psi_{1}$ and $\Psi_{2}$ to the ground-truth $\boldsymbol{\theta}^{*}$. $\Psi_{1}$ represents an asymptotic error bound of $\mathcal{O}\big{(}\frac{n}{\sqrt{m}}\big{)}$, while $\Psi_{2}$ represents an asymptotic error bound of $\mathcal{O}\big{(}\frac{n^{2}}{\sqrt{m}}\big{)}$. This indicates that, given the same amount of training data, our HPS achieves better preference alignment performance compared to PL. Specifically, the optimum solution $\boldsymbol{\theta}_{\text{HPS}}$ derived from HPS loss is closer to the desired ground-truth $\boldsymbol{\theta}^{*}$ than the solution obtained from PL loss. This suggests that HPS improves sample efficiency, making it advantageous in scenarios with limited data or when faster convergence to the true parameter is desired.
@@ 5.2 Reward Margin Analysis
Furthermore, to examine the global minimizer of our HPS training loss $\mathcal{L}_{{\boldsymbol{\theta}}}$, we analyze the optima of the training loss $\mathcal{L}_{\Pi}^{*}$ in Eqn. 12, since we have proved their good approximation in Theorem 2. Without loss of generality, when the number of dispreferred response samples $N=n-1\rightarrow\infty$, we can remove the $\log{N}$ from the HPS training loss $\mathcal{L}_{{\boldsymbol{\theta}}}$ as it does not change the minimizers and the geometry of the loss surface, and obtain a limiting objective:
@@
Table 3: Human evaluation comparing SFT, DPO-BT, DPO-PL, and DPO-HPS on user study dataset under fine-tuning conditions.
Method Quality Score
SFT 3.63
DPO-BT 3.82
DPO-PL 3.69
DPO-HPS 3.93
@@ 6 Experiments
Baselines. In our experiments, we employ a supervised fine-tuned Llama3-8B checkpoint RLHFlow/Llama3-SFT-v2.0 (Dong et al., 2024; Dubey et al., 2024) as both the naive baseline SFT and the reference model. In addition, we integrate three preference modeling strategies — BT, PL, and our proposed HPS — into several implicit reward parameterization frameworks, including DPO (Rafailov et al., 2024), EXO (Ji et al., 2024a), IPO (Azar et al., 2024), SPPO (Wu et al., 2024), and NCA (Chen et al., 2024). For example, DPO-PL refers to the configuration where Llama3-8B is fine-tuned using the DPO implicit reward parameterization under a PL preference model.
@@
Datasets. We use two popular datasets, HH-RLHF (Bai et al., 2022a) and PKU-SafeRLHF (Ji et al., 2024b), focusing on helpfulness and safety (Lambert et al., 2024; Fourrier et al., 2024). HH-RLHF is multi-turn, while PKU-SafeRLHF contains single question-answer pairs. Each prompt in the datasets includes two responses with human-rated preferences. Following prior work (Song et al., 2024), we expand response data by generating 100 responses using RLHFlow/Llama3-v2-DPO (Dong et al., 2024) per prompt. The corresponding rewards are computed via Skywork/Skywork-Reward-Llama3, a safety-aligned reward model (Liu et al., 2024a) ranked among the top 10 on the RewardBench (Lambert et al., 2024).
@@
Evaluation Metrics. We evaluate response quality and harmful content rejection. We use BLEU (Papineni et al., 2002) to assess the text quality by comparing responses to ground-truth preferred answers. To evaluate alignment with human preference, we also adopt a powerful reward model RLHFlow/ArmoRM-Llama3 (Wang et al., 2024), which is different from the one used during training, to measure the level of human preference gained. Importantly, we compute reward margins (RMs) from various implicit reward models (Table 1) to quantify the gap between preferred and dispreferred responses, where higher RM scores indicate better preference alignment without harmful or biased outputs.
@@
Human evaluation remains the gold standard for assessing response quality, where annotators compare two responses per question to select the better one or declare a tie. Thus, we conduct a user study to evaluate the performance of DPO and its variants. Moreover, recent LLMs like Qwen $2.5$ (Yang et al., 2024) closely align with human preferences, validated by benchmarks like the Open LLM Leaderboard (Fourrier et al., 2024), chatbot-arena-leaderboard (Chiang et al., 2024), and RewardBench (Lambert et al., 2024). We use Qwen2.5-72B-Instruct to assess response quality and win rates. For evaluation robustness, we sample $N=5$ responses per method and report the highest-scoring one for each metric.
@@
Implementation. Due to computational constraints, we apply LoRA (Hu et al., 2021) for efficient fine-tuning with a rank of $8$ and scaling factor $\alpha=16$. The KL penalty strength $\beta$ is set to $0.1$, following DPO. The ablation study about sensitivity of $\beta$ in DPO can be found in Appendix C. We fine-tune all methods for 2 epochs use AdamW optimizer (Loshchilov, 2017) with a learning rate of $5.0\times 10^{-7}$ over $2$ and a cosine learning rate scheduler. We set the sequence length as $2,048$ tokens for both training and inference, with a sampling temperature of $0.9$ during inference. In our HPS setting, we use the annotated scalar reward as the estimated reward $r_{\text{est}}$ for each response in Eqn. 9. The scaling factor $\gamma$ is scheduled to linearly increase from -5 to 5, with updates applied at every 20% interval of the training process. The GPU fine-tuning time for the PL-based, BT-based, and HPS-based methods is $168.4\!\pm\!1.9$ hours, $62.8\!\pm\!1.1$ hours, and $64.4\!\pm\!0.8$ hours, respectively. This demonstrates that our proposed HPS method can significantly improve efficiency compared to the PL-based method, achieving a $61.76\%$ reduction in fine-tuning time, and validates Theorem 1. More implementation details can be found in Appendix D.